Solve (x-9)(x+2)>0
The pypotenuse of a right triangle is 24 ft long. The length of one leg is 18 ft more than the other. Find the length of the legs. The length of the shorter leg is approximately ___________ft.
x^2+21<10x
f(x)=x^2-3x+1
The x intercept (s) is (are)_______
The y intercept (s) is (are)_______
Find the vertex, the line of symmetry, the maximum or minimum value of the quadratic function and graph the function: f(x)= -x^2+12x+2
Graph the function, label the vertex and draw the axis of symmetry:
f(x)=1/4(x-5)^2

Answer by feliz1965(151)

The hypotenuse of a right triangle is 24 ft long. The length of one leg is 18 ft more than the other. Find the length of the legs. The length of the shorter leg is approximately ___________ft.

We use the Pythagorean Theorem.
(leg)^2 + (leg)^2 = (hyp)^2
(x)^2 + (x + 18)^2 = (24)^2
x^2 + x^2 + 36x + 324 = 576
2x^2 + 36x + 324 - 576 = 0
2x^2 + 36x - 252 = 0
We can make this quadratic equation smaller by dividing every term by 2.
x^2 + 18x - 126 = 0
Can you finish now?